The ratio of orange juice to pinapple juice in tropical treat punch is 4 to 3 bill has 64 oz of orange juice how much pinapple juice does he need?

Answer by mananth(16946)

orange /pineapple = 4/3
4/3= 64/x
x=3*64/4
x=48 oz of pineapple juice
...
m.ananth@hotmail.ca